A sample of blood is placed in a centrifuge of radius 16.0 cm. The mass of a red blood cell is 3.0  10−16 kg, and the magnitude of the force acting on it as it settles out of the plasma is 4.0  10−11 N. At how many revolutions per second should the centrifuge be operated?


__rev/s


1
Expert's answer
2021-10-08T14:59:56-0400
"F=ma_c=m\dfrac{v^2}{r}=m\omega^2r,""\omega=\sqrt{\dfrac{F}{mr}},""\omega=\sqrt{\dfrac{4.0\cdot10^{-11}\ N}{3.0\cdot10^{-16}\ kg\cdot0.16\ m}}=913\ \dfrac{rad}{s},""\omega=913\ \dfrac{rad}{s}\cdot\dfrac{1\ rev}{2\pi\ rad}=145\ \dfrac{rev}{s}."

Answer:

"\omega=145\ \dfrac{rev}{s}."